OpenStudy (cookiemonster18):

Correct me if I'm wrong please 1. Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y = –2x + 4 –5x + 10y = 5 (1 point) A.parallel B.perpendicular * C.neither 2. Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y = -1/4x + 10 –2x + 8y = 6 (1 point) A.parallel B.perpendicular C.neither *

OpenStudy (anonymous):

is it clear that your first job is to identify the slope of each?

OpenStudy (cookiemonster18):

how do I identify the slope? @satellite73

OpenStudy (anonymous):

for \(y = –2x + 4 \) the slope is on display it is the coefficient of the x term

OpenStudy (anonymous):

for \[–5x + 10y = 5 \] you can solve for \(y\) to make it look like the previous one

OpenStudy (anonymous):

or just know that the slope of \[ax+by=c\] is \(-\frac{a}{b}\)

OpenStudy (anonymous):

you got the slope of \[y=-2x+4\]?

OpenStudy (anonymous):

ooh i am sorry, i see you have already picked an answer and just want a check

OpenStudy (cookiemonster18):

im very confused because I have no idea on how to solve this so am I suppose to put -2+4 into the calculator and find the answer??? im sorry im not good with this please help me solve these so i'd know how to correctly get the right answer

OpenStudy (anonymous):

first two are perpendicular like you said

OpenStudy (anonymous):

slope of \[y=mx+b\] is \(m\) so the slope of \(y=-2x+4\) is \(-2\)

OpenStudy (cookiemonster18):

so the first one is correct? ( perpendicular ) is the second one correct as well?

OpenStudy (anonymous):

slope of \[-5x+10y=5\] if you solve for y you get \[y=\frac{1}{2}x+\frac{1}{2}\] slope is \(\frac{1}{2}\)

OpenStudy (anonymous):

since \(\frac{1}{2}\) is the "negative reciprocal " of \(-2\) the lines are perpendicular

OpenStudy (cookiemonster18):

Okay thank you but what about the 2nd one?? can you show me how to solve that one as well @satellite73

OpenStudy (mathmale):

In Problem 2, start by identifying the SLOPES of the two given equations. What are they? Are the slopes equal? Are the slopes negative reciprocals of one another?

OpenStudy (cookiemonster18):

well since ig it's not neither then is it perpendicular?? @mathmale